# Synthesising Counterfactual Explanations via Label-Conditional Gaussian Mixture Variational Autoencoders

### v1

@@ 4.2 Evaluation Results
The benchmarking results are visualised in Figure 4. All included methods, apart from DiCE, always find valid CEs. In terms of proximity, NNCE and DRCE consistently perform the best, with DiCE showing competitive performances on continuous-only datasets. Our methods, especially LAPACE-First, show comparable performances. Notably, LAPACE achieves the best plausibility results across all datasets and models, effectively managing the known tradeoff between plausibility and proximity. In terms of diversity, our methods outperform DiCE and DRCE on datasets with mixed data types but perform worse on the two continuous-only datasets.
@@
Table 3 reports the actionability results of LAPACE against two FACE variants. We used a combination of constraints on 5 features per dataset, which is around 20% and 50% of total features on heloc and wine, respectively. Under these conditions, both LAPACE variants consistently found valid counterfactuals that satisfied all constraints, outperforming the baseline methods.
@@ 5 Conclusion
This work opens several exciting avenues for future research. Links to the synthetic tabular data generation literature (Stoian et al., 2025) are identified since it is how we view CE generation in this work. For example, dataset-level causal requirements between features (beyond user-specified actionability requirements) can be incorporated, possibly through adding specialised neural network layers into the training process of L-GMVAE (Stoian & Giunchiglia, 2025). Furthermore, the Gaussian mixture-regularised latent space learned by our model offers a powerful tool for extended forms of counterfactual explainability. Instead of generating point-based CEs for one input at a time, the structured manifold could be used to derive region-based explanations, providing insights into the model’s behaviour for entire sub-groups of the data (Rawal & Lakkaraju, 2020; Bewley et al., 2024).

### v2

@@ 4.2 Evaluation Results
Benchmarking results. The benchmarking results are visualised in Figure 4. All included methods, apart from DiCE, always find valid CEs. In terms of proximity, NNCE and DRCE consistently perform the best, with DiCE showing competitive performances on continuous-only datasets. Our methods, especially LAPACE-First, show comparable performances. Notably, LAPACE achieves the best plausibility results across all datasets and models, effectively managing the known tradeoff between plausibility and proximity. In terms of diversity, our methods outperform DiCE and DRCE on datasets with mixed data types but perform worse on the two continuous-only datasets.
@@
Actionability evaluation. Table 3 reports the actionability results of LAPACE against two FACE variants. We used a combination of constraints on 5 features per dataset, which is around 20% and 50% of total features on heloc and wine, respectively. Under these conditions, both LAPACE variants consistently found valid CEs that satisfied all constraints, outperforming the baselines.
Figure 5: Classifier-predicted probability of the CE target class and plausibility evaluation for all path points with varying $\tau$. Each subfigure in a row shows the results for a path obtained from each L-GMVAE cluster for the desirable class. Points with a class probability greater than or equal to 0.5 (grey dotted line) are classified as the desirable class and could be used as CE points.
More on the path points. To further examine the characteristics of all generated path points connecting the input to LAPACE-Last centroids, we evaluate the classifier’s predicted target class probabilities and the plausibility (LOF) of the paths. We present the results on heloc and wine in Figure 5. As $\tau$ increases, the class probabilities grow consistently, validating that the learned latent space of the L-GMVAE faithfully aligns with the classifier’s predictions. Note that the probabilities at smaller $\tau$ values are highly dependent on the random selection of test inputs; therefore, variance could be high there. LOF scores for entire paths, including those before crossing the decision boundary, are consistently low (near 1.0), indicating that these synthetic points lie well within the data manifold. The reconstructed cluster centroids (LAPACE-Last points) often exhibit the greatest plausibility with LOF below 1.0, confirming they are located in regions denser than their neighbours.
@@ 5 Conclusion
This work opens several exciting avenues for future research. Establishing theoretical robustness guarantees on CEs obtained leveraging L-GMVAE would be valuable (Jiang et al., 2024c). Links to the synthetic tabular data generation literature (Stoian et al., 2025) are identified since it is how we view CE generation in this work. For example, dataset-level causal requirements between features (beyond user-specified actionability requirements) can be incorporated, possibly through adding specialised neural network layers into the training process of L-GMVAE (Stoian and Giunchiglia, 2025). Furthermore, the Gaussian mixture-regularised latent space learned by our model offers a powerful tool for extended forms of counterfactual explainability. Advanced interpolation methods such as Bures-Wasserstein interpolation (Bhatia et al., 2019) could be tested for replacing our linear one. Instead of generating point-based CEs for one input at a time, the structured manifold could be used to derive region-based explanations, providing insights into the model’s behaviour for entire sub-groups of the data (Rawal and Lakkaraju, 2020; Bewley et al., 2024). Future work could also investigate extending L-GMVAE for synthesising CEs in other data modalities.
Acknowledgements
Jiang, Rago and Toni were partially funded by J.P. Morgan and by the Royal Academy of Engineering under the Research Chairs and Senior Research Fellowships scheme. Leofante was supported by an Imperial College Research Fellowship. Rago and Toni were partially funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101020934). Any views or opinions expressed herein are solely those of the authors listed.
@@ Appendix C L-GMVAE in the CE Evaluation
Appendix D The Use of Large Language Models
Large language model tools were used in a lightweight way to help structure the presentation of some paragraph drafts in the introduction and conclusion, at early stages of paper writing. They are not involved in further paper refinement. They are also used to aid in debugging the code implementations.